# FeatSharp: Your Vision Model Features, Sharper

### v1

@@ Appendix E Throughput Analysis
In (6), we predict that based on the quadratic scaling of attention, theoretically FeatSharp should always be cheaper than running the base model at the upsampled resolution. FeatSharp’s cost is linear in the number of tokens, whereas a ViT is quadratic. In figure 15, we show the results of this prediction on actual hardware. As can be seen with the “Actual” curve, the picture is a bit more complex than pure quadratic scaling, as between 1x and 3x upsample factors, the scaling is actually sub-linear, which likely reflects the period where self-attention is memory bound, and not compute bound, thus adding extra tokens doesn’t proportionally increase the cost. Specifically, at 1.85x upsampling, we achieve the lowest time per token, and from then on, the cost approximately linearly increases (note that time per token is the first derivative of the time per image, so linear growth implies quadratic scaling, as predicted). Because FeatUp only runs the featurizer once, and its upsampling operation is cheap, we can see that it achieves strong scaling regardless of resolution. FeatSharp requires $u^{2}+1$ inferences with $u$ being the upsample factor, so its cost is higher. Likely due to non-optimal kernels, we can see that FeatSharp does start operating faster than the base model until about 3.3x upsampling ($\approx 1260^{2}$px). However, also as predicted by (6), FeatSharp’s scaling is linear.
@@ Appendix F Effects of “Over-Tiling”
In figure 8, we can see that RADIO had learned some idiosyncratic representations when using the Tile and S2 upsampling algorithms. The effects are even more apparent in figures 9 and 10 where color spaces can entirely flip. To understand what’s happening, we rely on the pretrained RADIOv2.5-L model, which has strong scale equivariance properties (Heinrich et al., 2024), and first see that as the number of tiles increases, the MSE error between the brute-force inference at a given resolution and the tiling of that resolution, increases. We show these results in figure 16. Visually, we argue that the major increases in MSE owes largely to regions that lack context, making it difficult for the encoder (in this case RADIO), to come up with a reasonable representation of the tile-crop. We visualize this in figure 17. Notably, we can see that the $8\times 8$ tiling difference images are generally whiter, indicating a general drift towards higher error. We can also see particular tiles that have more error, such as the notecard in row 4, which gets nearly forgotten due to context. We can also see that there are a lot of errors with the car on row 5. The bottom center of the floor on row 6 has the same issue. So, while there appears to be a general upward error drift, it’s exacerbated in regions without much variation.

### v2

@@ Appendix E Throughput Analysis
E.1 Empirical Throughput
In (6), using $f(x)=c(1+x^{2})$ (e.g. non-progressive tiling), we predict that based on the quadratic scaling of attention, theoretically FeatSharp should always be cheaper than running the base model at the upsampled resolution. FeatSharp’s cost is linear in the number of tokens, whereas a ViT is quadratic. In figure 14, we show the results of this prediction on actual hardware. As can be seen with the “Actual” curve, the picture is a bit more complex than pure quadratic scaling, as between 1x and 3x upsample factors, the scaling is actually sub-linear, which likely reflects the period where self-attention is memory bound, and not compute bound, thus adding extra tokens doesn’t proportionally increase the cost. Specifically, at 1.85x upsampling, we achieve the lowest time per token, and from then on, the cost approximately linearly increases (note that time per token is the first derivative of the time per image, so linear growth implies quadratic scaling, as predicted). Because FeatUp only runs the featurizer once, and its upsampling operation is cheap, we can see that it achieves strong scaling regardless of resolution. FeatSharp requires $u^{2}+1$ inferences with $u$ being the upsample factor, so its cost is higher. Likely due to non-optimal kernels, we can see that FeatSharp does start operating faster than the base model until about 3.3x upsampling ($\approx 1260^{2}$px). However, also as predicted by (6), FeatSharp’s scaling is linear.
@@ E.1 Empirical Throughput
E.2 Proof of Equation 6
The progressive form of equation 6 is defined as $f(x)=\sum_{i=1}^{x}i^{2}$, and the regular form of self-attention is $g(x)=cx^{4}$, with $x$ being the upsampling factor per-side, and $c$ being the cost to evaluate at the base resolution. We want to show that:
$f(x)\leq g(x)\quad\forall x>1$ (11)
We start by rewriting the series for $f(x)$ in closed form
$f(x)=c\frac{x(x+1)(2x+1)}{6}$ (12)
which is the sum of squares sequence multiplied by $c$. So now
$f(x)\leq g(x)\Longleftrightarrow\frac{x(x+1)(2x+1)}{6}\leq x^{4}$ (13)
Given that $c>0$ and that it’s a constant factor on both sides, we can eliminate it.
$2x^{3}+3x^{2}+x\leq 6x^{4}$ (14)
and with $x>0$, we can further simplify to
$1\displaystyle 2x^{2}+3x+$ 1 $\leq 6x^{3}$ (15)
$1\displaystyle 6x^{3}-2x^{2}-3x-$ 1 $\geq 0$ (16)
$(x-1)(6x^{2}+4x+1)$ $\geq 0$ (17)
and thus $x-1\geq 0\quad\forall x\geq 1$, and also $6x^{2}+4x+1>0\quad\forall x\in\mathbb{R}$. Therefore, $f(x)\leq g(x)\quad\forall x>1$.
@@ Appendix F Effects of “Over-Tiling”
In figure 8, we can see that RADIO had learned some idiosyncratic representations when using the Tile and S2 upsampling algorithms. The effects are also apparent in figure 9 where color spaces can entirely flip. To understand what’s happening, we rely on the pretrained RADIOv2.5-L model, which has strong scale equivariance properties (Heinrich et al., 2024), and first see that as the number of tiles increases, the MSE error between the brute-force inference at a given resolution and the tiling of that resolution, increases. We show these results in figure 15. Visually, we argue that the major increases in MSE owes largely to regions that lack context, making it difficult for the encoder (in this case RADIO), to come up with a reasonable representation of the tile-crop. We visualize this in figure 16. Notably, we can see that the $8\times 8$ tiling difference images are generally whiter, indicating a general drift towards higher error. We can also see particular tiles that have more error, such as the notecard in row 4, which gets nearly forgotten due to context. We can also see that there are a lot of errors with the car on row 5. The bottom center of the floor on row 6 has the same issue. So, while there appears to be a general upward error drift, it’s exacerbated in regions without much variation.